OpenStudy (anonymous):

the edhelper florist sells roses at $11.72 a bunch, tulips at $4.50 a bunch, and daffodils at $5.22 a bunch. ryan ordered a banquet of 19 bunches of roses and tulips. if he wants to spend exactly $114.38, how many bunches of each can he buy?

OpenStudy (anonymous):

Let x = number of bunches of roses and y = number of bunches of tulips. From the information given, we can form the following equations: x + y = 19 (11.72)x + (4.5)y = 114.38 Solving the two equations will give you the answer.